Reviewing for finals. Getting stuck on these for some reason... Help is appreciated!

How many grams of NaCl would be needed to prepare a 2.0M NaCl solution? Molar mass of NaCl is 58.44g/mol

What is the molarity of a solution prepared by dissolving 3 mol NaI in 6L of solution?
Answer #1

molarity can be defined as = (no. of moles of solute)/(solution in liters)

here you need to know in what volume you want to prepare the solution
soppose you want to prepare a 2.0M in 1L

2=m/1L where m is no. of moles of solute
which when solved m=2moles
as 1mole NaCl = 58.44 g 2 moles will be 2*58.44g = 116.88g
it means if you want to prepare 2M solution in 1L you need 116.88g NaCl

let us suppose you want to prepare 2M solution in 0.5L

now
2=m/0.5L m=1mole
ie 58.44g
so if you want to prepare 0.5L 2molar solution you need 58.44g of NaCl

molarity of solution prepared by 3mole NaI in 6L =3/6=0.5M